# An Erdős problem on random subset sums in finite abelian groups

### v1

@@ 3.1. Reducing to an independent model
Recall that $M=2^{k}-1$ and $\lambda=M/p$. By Lemma 2.4, $\lambda=o(R)$. It suffices to establish the following uniform estimate
$\mathbb{E}[(X_{B})_{r}]=(m\lambda)^{r}\,(1+o(1))\qquad(1\leq r\leq R).$ (11)
As we will show later, once this estimate is proved, substituting it into (10) with $Z=X_{B}$ completes the proof.
To proceed, we define some notations. Write $a=(a_{1},\dots,a_{k})\in\mathbb{F}_{p}^{k}$. For each nonempty $S\subseteq[k]$, let $v(S)\in\{0,1\}^{k}$ be its indicator vector. So the definition (6) becomes $\sigma(S)=a\cdot v(S)$. Given subsets $S_{1},\dots,S_{r}\subseteq[k]$, we write $V=V(S_{1},\dots,S_{r})\in\{0,1\}^{r\times k}$ for the incidence matrix with entries $V_{j,t}=\mathbf{1}_{\{t\in S_{j}\}}$.
@@
$(X_{B})_{r}=\sum_{\begin{subarray}{c}(S_{1},b_{1}),\dots,(S_{r},b_{r})\\ \text{all distinct}\end{subarray}}\mathbf{1}_{\{\sigma(S_{1})=b_{1},\dots,\sigma(S_{r})=b_{r}\}}.$
@@
where $N_{r,d}(B)$ denotes the number of ordered $r$-tuples of distinct pairs $(S_{j},b_{j})$ for which the incidence matrix has rank $d$ (over $\mathbb{F}_{p}$) and the corresponding system is consistent. By Lemma 2.1, we may freely interpret ranks over $\mathbb{Q}$.
@@
Let $V$ be an incidence matrix of rank $d<r$ and let $W=\mathrm{Col}_{\mathbb{Q}}(V)\subseteq\mathbb{Q}^{r}$. The number of possible $W$ is at most the number of ordered $d$-tuples of vectors in $\{0,1\}^{r}$ spanning $W$, hence at most $(2^{r})^{d}\leq 2^{r^{2}}$. Now fix $W$. Every column of $V$ must lie in $\Gamma:=W\cap\{0,1\}^{r}$. By Lemma 2.2 and Lemma 2.3, we have $|\Gamma|\leq\frac{3}{4}\,2^{d}$. Thus the number of possible column sequences is at most $|\Gamma|^{k}$, giving the stated bound. ∎
@@
$\sum_{d<r}\frac{N_{r,d}(B)}{p^{d}}\leq m^{r}p^{-\eta_{0}+o(1)}\leq\frac{(mM)_{r}}{p^{r}}\,p^{-\eta}.$ (13)
For the full-rank case, note that when $\operatorname{rank}(V)=r$ the map $T$ (defined in the proof of Lemma 3.2) is surjective, so the system is consistent for every right-hand side. Thus the contribution from the full-rank case equals
$\frac{1}{p^{r}}\cdot\#\{(S_{1},b_{1}),\dots,(S_{r},b_{r}):\ \text{ all distinct pairs and }\operatorname{rank}(V)=r\}.$ (14)
To estimate (14), we need to show that almost all tuples of distinct pairs $(S_{1},b_{1}),...,(S_{r},b_{r})$ satisfy $\operatorname{rank}(V)=r$. For fixed $r\leq R$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}$, let $\mathcal{T}_{r}(B)$ be the set of ordered $r$-tuples of distinct pairs $(S_{i},b_{i})$ with $\varnothing\neq S_{i}\subseteq[k]$ and $b_{i}\in B$.
@@
Uniformly for all $1\leq r\leq R$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}$, the number of tuples in $\mathcal{T}_{r}(B)$ with $\mathrm{rank}(V)<r$ is at most $(mM)_{r}\,p^{-\eta_{1}}$ for some absolute constant $\eta_{1}>0$.
@@
We split the proof into two cases.
Case (i). Some subset repeats. This can only happen when $m=2$, and if $S_{i}=S_{j}$ then necessarily $b_{i}\neq b_{j}$ because the pairs are distinct. Fix $1\leq i<j\leq r$. There are at most $M$ choices for $S_{i}=S_{j}$, at most $2$ choices for $(b_{i},b_{j})\in B^{2}$ with $b_{i}\neq b_{j}$, and at most $(2M-2)_{r-2}\leq(2M)^{r-2}$ choices for the remaining pairs. Summing over $i<j$ gives
$\#\{\text{tuples with a repeated subset}\}\leq\binom{r}{2}\cdot M\cdot 2\cdot(2M-2)_{r-2}=\binom{r}{2}\cdot\frac{(2M)_{r}}{2M-1}\ll\frac{r^{2}}{M}\,(2M)_{r}.$
Since $r\leq R=(\log p)^{\beta}$ and $M=p^{1+o(1)}$, we have $r^{2}/M=p^{-1+o(1)}$, so this number is at most $(2M)_{r}\,p^{-1/2}$ for all large $p$.
Case (ii). All subsets are distinct. For each ordered $r$-tuple of distinct nonempty subsets $(S_{1},\dots,S_{r})$ with $\mathrm{rank}(V)=d<r$, there are at most $m^{r}$ choices of $(b_{1},\dots,b_{r})\in B^{r}$. Recall that $T_{r,d}$ is the number of ordered $r$-tuples of distinct nonempty subsets $(S_{1},\dots,S_{r})$ of $[k]$ whose incidence matrix $V$ has rank $d$. Hence
$\{\text{tuples with all }S_{i}\text{ distinct and }\mathrm{rank}(V)=d\}\ \leq\ m^{r}\,T_{r,d}.$
By Lemma 3.3,
$\sum_{d<r}m^{r}T_{r,d}\ \leq\ m^{r}\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\sum_{d<r}(2^{k})^{d}\ \leq\ m^{r}\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\cdot r\,(2^{k})^{r-1}.$
Using $M=2^{k}-1\asymp 2^{k}$ and (2), we know that $(mM)_{r}\asymp(mM)^{r}\asymp m^{r}(2^{k})^{r}$ uniformly for $r\leq R$. Let $\eta_{2}:=-\log_{2}(3/8)>0$. Putting the above bounds together, for all large $p$ we have
$\frac{\#}{(mM)_{r}}\leq\ r\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\cdot\frac{(2^{k})^{r-1}}{(2^{k})^{r}}\ =\ r\,2^{r^{2}}\Bigl(\tfrac{3}{8}\Bigr)^{k}\ =\ p^{-\eta_{2}+o(1)}\leq p^{-\eta_{2}/2}.$
Combining both cases, we conclude that the lemma holds with $\eta_{1}:=\min\{1/2,\eta_{2}/2\}$. ∎
Fix $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}.$ Since $|\mathcal{T}_{r}(B)|=(mM)_{r}$, by (14) and Lemma 3.4, the contribution of the full-rank case in (12) equals
$\frac{\#\{(S_{i},b_{i})\in\mathcal{T}_{r}(B):\ \mathrm{rank}(V)=r\}}{p^{r}}=\frac{(mM)_{r}-O\mathopen{}\mathclose{{\left((mM)_{r}p^{-\eta_{1}}}}\right)}{p^{r}}=\frac{(mM)_{r}}{p^{r}}\,(1+O(p^{{-\eta_{1}}})),$
uniformly for all $1\leq r\leq R$. Together with (12) and (13) this yields
$\mathbb{E}[(X_{B})_{r}]=\frac{(mM)_{r}}{p^{r}}\bigl(1+O(p^{-\eta_{3}})\bigr)\qquad(1\leq r\leq R),$
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|\in\{1,2\}$, where $\eta_{3}:=\min\{\eta/2,\eta_{1}/2\}>0$. By (2), we have $(mM)_{r}=(mM)^{r}(1+o(1))$. Since $M/p=\lambda$, we derive that
$\mathbb{E}[(X_{B})_{r}]=(m\lambda)^{r}\bigl(1+O(p^{-\eta_{3}})\bigr)\qquad(r\leq R),$ (15)
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|\in\{1,2\}$, proving (11).
@@
$\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{\mathbb{E}[(X_{B})_{r}]}{r!}=\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{(m\lambda)^{r}}{r!}+O\mathopen{}\mathclose{{\left(p^{-\eta_{3}}\sum_{r=0}^{\widetilde{R}}\frac{(m\lambda)^{r}}{r!}}}\right),$
@@
Using $\sum_{r=0}^{\widetilde{R}}(m\lambda)^{r}/r!\leq e^{m\lambda}$ and $e^{m\lambda}=p^{o(1)}$ (by Lemma 2.4), the error $O(p^{-\eta_{3}}e^{m\lambda})$ is $o(e^{-m\lambda})$. Moreover, since $\lambda=o(R)$, Stirling’s formula yields
@@ 4. Concluding Remarks
This demonstrates that the upper bound $f(N)\leq\log_{2}N+o(\log\log N)$ cannot hold uniformly over all finite abelian groups of order $N$. Nevertheless, certain abelian groups $G$ of order $N$ do exist for which, with high probability, a uniformly random subset of size $\log_{2}N+o(\log\log N)$ (or even smaller size) generates all elements of $G$. For such examples, we refer to Sothanaphan [12], which treats the case $G=(\mathbb{Z}/2\mathbb{Z})^{d}$.
Erdős and Rényi [8, Theorem 2] proved that for all $N$,
@@
Determine the following constant $c_{*}$ satisfying
$c_{*}=\limsup_{N\to\infty}\frac{f(N)-\log_{2}N}{\log\log N}$
Using (16) and (17), we can derive that $\frac{1}{2\log 2}\leq c_{*}\leq\frac{1}{\log 2}$. Heuristics and numerical experiments suggest that the constant in Problem 4.1 is very likely $c_{*}=\tfrac{1}{\log 2}$. However, the present method does not reach the regime $c_{*}>\tfrac{1}{2\log 2}$. One evident bottleneck is the proof of Lemma 3.3. It would be very interesting to push the lower bound constant beyond $\tfrac{1}{2\log 2}$.
Acknowledgements
We thank Wouter van Doorn for helpful suggestions that improved the presentation of this paper. We are grateful to Nat Sothanaphan for providing relevant materials in [12] and for an elegant rewrite [11] of the original AI-generated proof. We also thank Boris Alexeev, Ingo Althöfer, Thomas Bloom, Mehtaab Sawhney, Terence Tao, and Shengtong Zhang for helpful discussions on the Erdős Problems website forum. Finally, we are grateful to Thomas Bloom for founding and maintaining the Erdős Problems website [4]. This work is supported by National Key Research and Development Program of China 2023YFA1010201, National Natural Science Foundation of China grant 12125106, and Innovation Program for Quantum Science and Technology 2021ZD0302902.

### v2

@@ 3.1. Reducing to an independent model
Recall that $M=2^{k}-1$ and $\lambda=M/p$. Since $\lambda=o(R)$, it is enough for the Bonferroni step to prove that there exists a constant $\eta_{*}>0$ [^1] Later in the proof, we will rename this constant as $\eta_{4}$ once all error terms have been assembled. such that
$\mathbb{E}[(X_{B})_{r}]=(m\lambda)^{r}\bigl(1+O(p^{-\eta_{*}})\bigr)\qquad(1\leq r\leq R),$ (10)
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m$. As we will show later, once this estimate is proved, substituting it into (9) with $Z=X_{B}$ completes the proof.
We next introduce some notation. Write $a=(a_{1},\dots,a_{k})\in\mathbb{F}_{p}^{k}$. For each nonempty $S\subseteq[k]$, let $v(S)\in\{0,1\}^{k}$ be its indicator vector. So the definition (7) becomes $\sigma(S)=a\cdot v(S)$. Given subsets $S_{1},\dots,S_{r}\subseteq[k]$, we write $V=V(S_{1},\dots,S_{r})\in\{0,1\}^{r\times k}$ for the incidence matrix with entries $V_{j,t}=\mathbf{1}_{\{t\in S_{j}\}}$.
@@
$(X_{B})_{r}=\sum_{\begin{subarray}{c}(S_{1},b_{1}),\dots,(S_{r},b_{r})\\ \text{all distinct}\end{subarray}}\mathbf{1}_{\{\sigma(S_{1})=b_{1},\dots,\sigma(S_{r})=b_{r}\}}=\sum_{\begin{subarray}{c}S_{1},\dots,S_{r}\subseteq[k]\\ \varnothing\neq S_{i},\ \text{all distinct}\end{subarray}}\ \sum_{b_{1},\dots,b_{r}\in B}\mathbf{1}_{\{\sigma(S_{1})=b_{1},\dots,\sigma(S_{r})=b_{r}\}}.$
Here the second equality is valid because if $S_{i}=S_{j}$ and $b_{i}\neq b_{j}$, then the corresponding indicator is identically zero, while if $S_{i}=S_{j}$ and $b_{i}=b_{j}$, the pairs $(S_{i},b_{i})$ and $(S_{j},b_{j})$ are not distinct.
@@
where $N_{r,d}(B)$ denotes the number of ordered tuples
$(S_{1},\dots,S_{r};\,b_{1},\dots,b_{r}),$
with $S_{1},\dots,S_{r}$ distinct nonempty subsets of $[k]$ and $b_{1},\dots,b_{r}\in B$, for which the incidence matrix has rank $d$ over $\mathbb{F}_{p}$ and the corresponding system is consistent. By Lemma 2.1, we may freely interpret ranks over $\mathbb{Q}$.
@@
Let $V$ be an incidence matrix of rank $d<r$ and let $W$ denote the column space of $V$ over $\mathbb{Q}$. The number of possible $W$ is at most the number of ordered $d$-tuples of vectors in $\{0,1\}^{r}$ spanning $W$, hence at most $(2^{r})^{d}\leq 2^{r^{2}}$. Now fix $W$. Every column of $V$ must lie in $\Gamma:=W\cap\{0,1\}^{r}$. By Lemma 2.2 and Lemma 2.3, we have $|\Gamma|\leq\frac{3}{4}\,2^{d}$. Thus the number of possible column sequences is at most $|\Gamma|^{k}$, giving the stated bound. ∎
@@
$\sum_{d<r}\frac{N_{r,d}(B)}{p^{d}}\leq rm^{r}p^{-\eta_{0}+o(1)}\leq\frac{(mM)_{r}}{p^{r}}\,p^{-\eta}.$ (12)
For the full-rank case, note that when $\operatorname{rank}(V)=r$, the map $T$ (defined in the proof of Lemma 3.2) is surjective, so the system is consistent for every right-hand side. Thus the contribution from the full-rank case equals
$\frac{m^{r}}{p^{r}}\cdot\#\Bigl\{(S_{1},\dots,S_{r}):\ \varnothing\neq S_{i}\subseteq[k],\ \text{all distinct},\ \operatorname{rank}(V)=r\Bigr\}.$ (13)
To estimate (13), it suffices to show that almost all ordered $r$-tuples of distinct nonempty subsets of $[k]$ have full rank.
@@
Uniformly for all $1\leq r\leq R$, the number of ordered $r$-tuples of distinct nonempty subsets $(S_{1},\dots,S_{r})$ of $[k]$ with $\operatorname{rank}(V)<r$ is at most
$(M)_{r}\,p^{-\eta_{1}}$
for some absolute constant $\eta_{1}>0$.
@@
Let $T_{r,d}$ be as in Lemma 3.3. By the same argument as in the proof of Lemma 3.3, but using only the trivial bound
$|W\cap\{0,1\}^{r}|\leq 2^{d},$
we obtain
$T_{r,d}\leq 2^{r^{2}}(2^{d})^{k}\qquad(d<r).$
Hence
$\{(S_{1},\dots,S_{r}):\ \text{all distinct and }\operatorname{rank}(V)<r\}=\sum_{d<r}T_{r,d}\leq 2^{r^{2}}\sum_{d<r}(2^{d})^{k}\leq r\,2^{r^{2}}(2^{k})^{r-1}.$
On the other hand, since $r\leq R$ and $M=2^{k}-1\asymp 2^{k}$, we have
$(M)_{r}\asymp(2^{k})^{r}$
uniformly for $r\leq R$ (for instance by (3) with $m=1$). Therefore
$\frac{\#\{(S_{1},\dots,S_{r}):\ \text{all distinct and }\operatorname{rank}(V)<r\}}{(M)_{r}}\ll r\,2^{r^{2}}\,2^{-k}=p^{-1+o(1)}.$
Since $r\leq R\leq(\log p)^{\beta}$, we have $2^{r^{2}}=p^{o(1)}$, while $2^{-k}=p^{-1+o(1)}$. Thus the last quantity is at most $p^{-1/2}$ for all sufficiently large $p$. This proves the lemma. ∎
By Lemma 3.4, the quantity in (13) equals
$\frac{m^{r}}{p^{r}}\Bigl((M)_{r}+O\bigl((M)_{r}p^{-\eta_{1}}\bigr)\Bigr)=\frac{m^{r}(M)_{r}}{p^{r}}(1+O(p^{-\eta_{1}})).$
Since (3) (applied first with $m=1$, and then with the fixed value $m\in\{1,2\}$) gives
$(M)_{r}=M^{r}\exp\Bigl(O\Bigl(\frac{r^{2}}{M}\Bigr)\Bigr)\qquad\text{and}\qquad(mM)_{r}=(mM)^{r}\exp\Bigl(O\Bigl(\frac{r^{2}}{M}\Bigr)\Bigr)$
uniformly for $r\leq R$, it follows that
$m^{r}(M)_{r}=(mM)_{r}\exp\Bigl(O\Bigl(\frac{r^{2}}{M}\Bigr)\Bigr).$
Now $r\leq R=(\log p)^{\beta}$ and $M=2^{k}-1\asymp p(\log p)^{\alpha}$, so $r^{2}/M=p^{-1+o(1)}$. Hence there exists an absolute constant $\eta_{1}^{\prime}>0$ such that $m^{r}(M)_{r}=(mM)_{r}\bigl(1+O(p^{-\eta_{1}^{\prime}})\bigr)$ uniformly for $r\leq R$, and so the contribution from the full-rank case is
$\frac{(mM)_{r}}{p^{r}}\bigl(1+O(p^{-\eta_{1}^{\prime\prime}})\bigr),$
uniformly for all $1\leq r\leq R$, where $\eta_{1}^{\prime\prime}:=\min\{\eta_{1},\eta_{1}^{\prime}\}>0$. Together with (11) and (12) this yields
$\mathbb{E}[(X_{B})_{r}]=\frac{(mM)_{r}}{p^{r}}\bigl(1+O(p^{-\eta_{2}})\bigr)\qquad(1\leq r\leq R),$
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|\in\{1,2\}$, where $\eta_{2}:=\min\{\eta/2,\eta_{1}^{\prime\prime}/2\}>0$. By (3), we have
$(mM)_{r}=(mM)^{r}\exp\Bigl(O\Bigl(\frac{r^{2}}{M}\Bigr)\Bigr)$
uniformly for $r\leq R$. Since $r^{2}/M=p^{-1+o(1)}$, there exists an absolute constant $\eta_{3}>0$ such that $(mM)_{r}=(mM)^{r}\bigl(1+O(p^{-\eta_{3}})\bigr)$ uniformly for $r\leq R$. As $M/p=\lambda$, it follows that
$\frac{(mM)_{r}}{p^{r}}=(m\lambda)^{r}\bigl(1+O(p^{-\eta_{3}})\bigr).$
Therefore
$\mathbb{E}[(X_{B})_{r}]=(m\lambda)^{r}\bigl(1+O(p^{-\eta_{4}})\bigr)\qquad(r\leq R),$ (14)
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|\in\{1,2\}$, where $\eta_{4}:=\min\{\eta_{2},\eta_{3}\}>0$, proving (10).
@@
$\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{\mathbb{E}[(X_{B})_{r}]}{r!}=\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{(m\lambda)^{r}}{r!}+O\mathopen{}\mathclose{{\left(p^{-\eta_{4}}\sum_{r=0}^{\widetilde{R}}\frac{(m\lambda)^{r}}{r!}}}\right),$
@@
Using $\sum_{r=0}^{\widetilde{R}}(m\lambda)^{r}/r!\leq e^{m\lambda}$ and $e^{m\lambda}=p^{o(1)}$ (by Lemma 2.4), the error $O(p^{-\eta_{4}}e^{m\lambda})$ is $o(e^{-m\lambda})$. Moreover, since $\lambda=o(R)$, Stirling’s formula yields
@@ 4. Concluding Remarks
This demonstrates that the upper bound $f(N)\leq\log_{2}N+o(\log\log N)$ cannot hold uniformly over all finite abelian groups of order $N$. Nevertheless, certain abelian groups $G$ of order $N$ do exist for which, with high probability, a uniformly random subset of size $\log_{2}N+o(\log\log N)$ (or even smaller size) generates all elements of $G$. For a related discussion in the special case $G=(\mathbb{Z}/2\mathbb{Z})^{d}$, see Sothanaphan [10].
Erdős and Rényi [6, Theorem 2] proved that for all $N$,
@@
Determine the value of
$c_{*}:=\limsup_{N\to\infty}\frac{f(N)-\log_{2}N}{\log\log N}.$
Using (15) and (16), we can derive that $\frac{1}{2\log 2}\leq c_{*}\leq\frac{1}{\log 2}$. It is plausible that the upper bound is sharp, but the present method does not reach the regime $c_{*}>\frac{1}{2\log 2}$. One evident bottleneck is the proof of Lemma 3.3. It would be very interesting to push the lower bound constant beyond $\frac{1}{2\log 2}$.
Acknowledgements. We thank Wouter van Doorn for helpful suggestions that improved the presentation of this paper. We are grateful to Nat Sothanaphan for providing relevant materials in [10]. We also thank Boris Alexeev, Ingo Althöfer, Thomas Bloom, Mehtaab Sawhney, Terence Tao, and Shengtong Zhang for helpful discussions on the Erdős Problems website forum. Finally, we are grateful to Thomas Bloom for founding and maintaining the Erdős Problems website [2]. This work is supported by National Key Research and Development Program of China 2023YFA1010201, National Natural Science Foundation of China grant 12125106, and Innovation Program for Quantum Science and Technology 2021ZD0302902.
Disclosure. An AI assistant was used as an exploratory aid at an early stage of this project. All statements, proofs, and the final presentation in this paper were independently verified and written by the authors. A fuller historical note on the exploratory phase is recorded in the arXiv version of this paper [9, Section 1.1].